# Christian Peukert [^1] Corresponding author. Florian Abeillon Jérémie Haese Franziska Kaiser Alexander Staub

### v1

@@ 4.5 Identification strategy and econometric model
Our causal identification strategy is based on the notion that the images in the LITE dataset are randomly drawn from a superset of all images on the platform that meet the eligibility criteria. This implies quasi-randomization at the user level. For our econometric analysis, we rely on a difference-in-differences approach. We perform a set of OLS regressions saturated with fixed effects to compare users with images in the LITE dataset to users who had at least one image deemed eligible to be included in the LITE dataset but were not included in the LITE dataset. Our baseline specification is
@@
where $Y_{it}$ is the outcome of interest, e.g. the number of images uploaded by user $i$ in month $t$. $Treated_{i}$ indicates whether user $i$ was included in the LITE dataset. $Post_{t}$ is a binary variable that takes the value 1 if the current month is after the release of the LITE dataset, and 0 otherwise. We include month fixed effects $\eta_{t}$ in all specifications, and in some specifications, we additionally include user fixed effects $\mu_{i}$. Because inclusion in the LITE dataset is the result of a natural experiment with randomization, the parameter $\delta$ estimates the causal effect of being included in a dataset made available for AI training on user behavior.
@@ 5 Results
We begin by reporting baseline results separately for users’ deletion rates as well as their uploading behavior. Then, we discuss effect heterogeneity by different types of users and different types of images.
@@ 5.1.1 Changes in deletion rates
First, we study whether being included in the LITE dataset affects the likelihood of an image or a user account being deleted from the platform. Hence, this set of results differs from all other analyses else below, which focus on users’ uploading behavior.
@@
Note: We estimate a linear probability model. The dependent variable is an indicator of whether an image or a user account is still available on Unsplash in May 2023. Treated indicates whether a user was part of the LITE dataset. We compare the survival of (images by) users who were part of the LITE dataset to the survival of (images by) users who were not part of LITE dataset over a period of 2.5 years (from Aug 20 to May 23). Fixed effects for the image’s age (in months). Standard errors clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
In column (1) of Table 2, we see that images that were included in the LITE dataset (i.e. that were treated) have a 2 percentage points higher chance of surviving on the platform than images that were not included in the LITE dataset. In column (2), we limit the sample to images of photographers that have at least two uploads and find the same result. We do so to compare against the specification in column (3), which has user-fixed effects. The coefficient there implies a 1 percentage point higher likelihood of survival. This suggests substantial differences in the variation between users and within users. Overall the results suggest that treated images are deleted at a rate of 7-8% ($\approx$ 3% per year), whereas control images are deleted at a rate of 9% ($\approx$ 4% per year). In columns (4) and (5), we turn to user accounts. We see that treated user accounts are less likely to survive than accounts of control users. The estimate is about 1 percentage point in both samples of all users and those users with at least two observations (i.e., the users in the sample in columns 2 and 3). This implies that accounts of treated users are deleted at a rate of 3% ($\approx$ 1% per year), whereas control users delete their accounts at a rate of 4% ($\approx$ 1.5% per year).
@@ 5.1.2 Changes in users’ uploading behavior
Panel A: Levels
Panel B: Differences
Note: In Panel A, we depict the average number of uploaded images for users in the treatment group (solid) and users in the control group (dashed). In Panel B, we plot OLS estimates of the $\delta_{\tau}$ coefficients are obtained from $Y_{it}=\sum_{\tau\in T}\nu_{\tau}+\sum_{\tau\in T}\delta_{\tau}\left(\gamma_{\tau}\times Treated_{i}\right)+\mu_{i}+\varepsilon_{it}$, with the number of uploads of user $i$ per month $t$ as the dependent variable. The dots reflect month-specific point estimates comparing treated users to the control group. Standard errors are clustered at the user level, and bars indicate 90% confidence bands. In both panels, the vertical line indicates the creation of LITE 1.0.0.
In the following specifications, we focus on the number of new uploads as our dependent variable. We compare users that had at least one image in the LITE dataset to our control group. For this analysis, we can draw upon high-frequency information, which allows us to construct a panel dataset at the user-month level. In Panel A of Figure 1, we plot the average number of uploads per month in the treatment group and the control group. In the observed 18 months before the creation of the LITE dataset (indicated by a 0 on the horizontal axis), upload rates do not fluctuate much for both. Treated users upload on average around 3 images per month and control users upload around 1 image per month. Looking at differences between treated and control in Panel B of Figure 1 confirms that there are no statistically significant differences in the before period. However, we see a substantial and statistically significant decrease in the after period. The effect is immediate and remains relatively stable over time. After 24 months, there seems to be a further decrease, which we investigate further in section 5.4.
@@
Post $\times$ Treated -1.1099∗∗∗ -0.0691∗∗∗ -0.1630∗∗∗
(0.09693) (0.00352) (0.00728)
Mean Treated Before 2.9038 0.2635 0.4892
Month FE Yes Yes Yes
User FE Yes Yes Yes
Observations 612,662 612,662 612,662
@@
We can quantify the average treatment effect for the entire observed post-period in an OLS model specified as described in equation (1). The results are reported in Table 3. In column (1), we look at the number of uploaded photos per month and quantify the average treatment effect as 1 less upload per month, which is a reduction of about 38%. We get similar results when we apply a log transformation in column (3). Further, in column (2), we document a decrease in the extensive margin of about 30%, i.e., when we use a dependent variable that indicates whether a user has uploaded at least one image in a given month.
5.2 Changes by different types of users
5.2.1 Professional versus amateur photographers
Beyond the average treatment effect on the treated, investigating heterogeneity in user characteristics can provide further nuance. Hence, we consider whether behavior differs between professional and amateur users. Thereafter, we turn to differences between more or less prolific users.
@@ 5.2.1 Professional versus amateur photographers
Note: The dependent variable is the number of uploads in a given month. Treated indicates whether a user was part of the LITE dataset. Pro-Gear indicates whether the user has uploaded at least one photo taken with professional gear (see section 4 for a detailed definition). ForHire indicates whether a user has chosen to show “Available for hire” on their Unsplash profile page. The number of observations differs from those in Table 3 because we do not observe Pro-Gear for all images and ForHire for all users. Fixed effects for upload month of the image and user account. Standard errors clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$ $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
The results in column (1) of Table 4 show that treated users who upload photographs created with professional gear react even more negatively than others. Similarly, we find in column (2) that treated users with a badge on their profile page that says “Available for hire” reduced their uploads disproportionally compared to users without a badge. It is possible that users with professional gear and those who signal that they are available for hire make a living as photographers and are, therefore, perhaps more directly affected by the potential displacement effects of AI. Hence, monetary motives may be one of the drivers for users to contribute less to the platform after their works have been made available for AI research.
5.2.2 Heterogeneity in contribution intensity
@@
The distinction between professional and amateur photographers is helpful to understand the societal impact of AI. Another interesting dimension of user heterogeneity concerns upload activity, which is an important metric to draw insights for platform governance. In Table 5, we show that the main effect reported in column (1) of Table 3 varies by percentile of a user’s total upload activity. The point estimates in columns (1)–(4) of Table 5 are smaller and not always significantly different from the estimate in column (1) of Table 3 in the sense that 90% confidence bands overlap. However, in absolute terms, the implied effect sizes are larger (column 2: -45%, column 3: -44% and column 4: -40%) compared to the -30% of the baseline. Given that we observe the universe of activity on Unsplash, we prefer not to remove particularly prolific users and report estimates of average treatment effects based on all users as our more conservative baseline results.
Table 5: Results: Changes in uploads, by user activity
(1) (2) (3) (4)
All w/o 99th w/o 95th w/o 90th
Post $\times$ Treated -0.4529∗∗∗ -0.9872∗∗∗ -0.6847∗∗∗ -0.4571∗∗∗
(0.02672) (0.05445) (0.03513) (0.02699)
Post $\times$ Treated $\times$ Above99thPct -17.4338
Post $\times$ Treated $\times$ Above95thPct -0.9941
Post $\times$ Treated $\times$ Above90thPct -1.3862
Mean Treated Before 2.9038 2.2160 1.5488 1.1435
Month FE Yes Yes Yes Yes
User FE Yes Yes Yes Yes
Observations 612,662 606,490 581,988 551,076
5.3 Changes in the types of images
5.3.1 Nature-themed and curated images

### v2

@@ 3.2.3 Identification strategy and econometric model
For our econometric analysis, we rely on a difference-in-differences approach. We perform a set of OLS regressions saturated with fixed effects to compare users with images in the LITE dataset to users who had at least one image deemed eligible to be included in the LITE dataset but were not included in the LITE dataset. Our baseline specification is
@@
where $Y_{it}$ is the outcome of interest, e.g., the number of images uploaded by user $i$ in month $t$. $Treated_{i}$ indicates whether user $i$’s works were included in the LITE dataset. $Post_{t}$ is a binary variable that takes the value 1 if the current month is after the release of the LITE dataset, and 0 otherwise. We include month fixed effects $\eta_{t}$ in all specifications, and in some specifications, we additionally include user fixed effects $\mu_{i}$. User fixed effects capture any unobserved time-invariant variation across users, such as the higher likelihood of prolific contributors having their images included in the LITE dataset. The parameter $\delta$ estimates the causal effect of being included in a dataset made available for AI training on user behavior.
@@ 4 Results
We begin by reporting baseline results separately for users’ deletion rates as well as their uploading behavior. Then, we explore potential underlying mechanisms.
@@ 4.1.1 Changes in deletion rates
First, we study whether being included in the LITE dataset affects the likelihood of an image or a user account being deleted from the platform. Hence, this set of results differs from all other analyses below, which focus on users’ uploading behavior.
@@
Note: We estimate a linear probability model. The dependent variable is an indicator for whether an image (columns 1-3) or a user account (columns 4-5) remains on Unsplash as of May 2023. In column (1), the sample includes all images. In column (2), the sample is restricted to images of user accounts with at least two images in the LITE dataset. In column (3), we add user fixed effects. In columns (4) and (5), the outcome switches to the user level (i.e., whether a user account is still active). Column (4) includes all users, and column (5) again restricts to users with at least two images in the LITE dataset. The variable Treated indicates whether a user was part of the LITE dataset. We compare survival rates over 2.5 years (from Aug 2020 to May 2023). We include fixed effects for the image’s age (in months) in all specifications. Standard errors are clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
In column (1) of Table 2, we see that images that were included in the LITE dataset (i.e., that were treated) have a 2 percentage points higher chance of surviving on the platform than images that were not included in the LITE dataset. [^19] Calculated as the coefficient (column 1: 0.0220) divided by the mean survival rate of the control group (columns 1-3: 0.9075) In column (2), we limit the sample to works of photographers that have at least two images included in the LITE dataset and find the same result. We do so to compare against the specification in column (3), which has user-fixed effects. The coefficient there implies a 1 percentage point higher likelihood of survival. This suggests substantial differences in the variation between users and within users. Overall the results suggest that treated images are deleted at a rate of 7-8% ($\approx$ 3% per year), whereas control images are deleted at a rate of 9% ($\approx$ 4% per year) [^20] For the control group, the deletion rate is calculated as 1 minus the mean survival rate of the control group (columns 1-3: 0.9075) while the coefficient (column 1: 0.0220) is added to the control group’s mean survival rate before subtracting to arrive at the treated group’s deletion rate. The yearly approximate deletion rate is derived by averaging across the time available in our sample (2.5 years).. The observed effects on deletion likely stem from substantial heterogeneity in user responses, consistent with anecdotal evidence we discuss in Section 4.2. We provide quantitative evidence supporting this heterogeneity in Table A.1. Specifically, users investing in expensive camera gear are significantly less likely to maintain their images on the platform (significant in column 3), aligning with qualitative evidence suggesting professional photographers are particularly critical of their data use for AI. Conversely, images belonging to users who subsequently display a ”For Hire” show increased persistence on Unsplash, indicating perceived business opportunities outweigh potential AI displacement concerns. Furthermore, images from users experiencing more intense treatment (those with multiple images included in the LITE dataset) exhibit higher deletion likelihood, supporting the argument presented later (Section 4.2.3) that more intensely treated users significantly reduce their future upload activity.
In columns (4) and (5), we turn to user accounts. We see that treated user accounts are less likely to survive than accounts of control users. The estimate is about 1 percentage point in both samples of all users and those users with at least two observations (i.e., the users in the sample in columns 2 and 3). This implies that accounts of treated users are deleted at a rate of 3% ($\approx$ 1% per year), whereas control users delete their accounts at a rate of 4% ($\approx$ 1.5% per year).
@@ 4.1.2 Changes in users’ uploading behavior
A: Levels
B: Differences
Note: In Panel A, we depict the demeaned average number of uploaded images for users in the treatment group (solid) and users in the control group (dashed). We use the group-specific pre-period mean number of uploaded images to demean (3.44 for the treatment group and 1.33 for the control group). In Panel B, we plot OLS estimates of the $\delta_{\tau}$ coefficients obtained from estimating $Y_{it}=\sum_{\tau\in T}\nu_{\tau}+\sum_{\tau\in T}\delta_{\tau}\left(\gamma_{\tau}\times Treated_{i}\right)+\mu_{i}+\varepsilon_{it}$, with the number of uploads of user $i$ per month $t$ as the dependent variable. The dots reflect month-specific point estimates comparing treated users to the control group. Standard errors are clustered at the user level, and bars indicate 90% confidence bands. In both panels, the vertical line indicates the creation of LITE 1.0.0.
We now focus on the number of new uploads as the dependent variable. We compare users with at least one image in the LITE dataset to the control group. For this analysis, we can draw upon high-frequency information, which allows us to construct a panel dataset at the user-month level. In Panel A of Figure 1, we plot the average number of uploads per month in the treatment group and the control group. For ease of interpretation, we demean using group-specific pre-period values. In the observed 18 months before the creation of the LITE dataset (indicated by zero on the horizontal axis), upload rates do not fluctuate much for both. Looking at differences between treated and control in Panel B of Figure 1 confirms that there are no statistically significant differences in the before period (all month-specific differences are jointly not significantly different from zero). The lack of statistically significant deviations from 0 in the pre-period supports the parallel trends assumption of the difference-in-differences model. After the policy change is communicated, we see a substantial and statistically significant decrease in the after period. The effect is immediate and remains relatively stable over time. After 24 months, there seems to be a further decrease, which we investigate further in section 4.2.3.
@@
(1) (2) (3) (4) (5) (6)
Post $\times$ Treated -1.0623∗∗∗ -1.1099∗∗∗ -0.0694∗∗∗ -0.0691∗∗∗ -0.1605∗∗∗ -0.1630∗∗∗
(0.09318) (0.09693) (0.00354) (0.00352) (0.00717) (0.00728)
Mean Treated Before 2.9038 2.9038 0.2635 0.2635 0.4892 0.4892
Month FE Yes Yes Yes Yes Yes Yes
User FE No Yes No Yes No Yes
Observations 612,662 612,662 612,662 612,662 612,662 612,662
Note: The dependent variable in columns (1) and (2) is the total number of uploads per month, in columns (3) and (4) it is an indicator of whether a user account has uploaded at least one image in a given month, and it is the log number of uploads in columns (5) and (6). Treated indicates whether a user was part of the LITE dataset. All specifications include fixed effects for the upload month of the image; the specifications in columns (2), (4), and (6) additionally include user-fixed effects. Standard errors are clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
@@
We can quantify the average treatment effect for the entire observed post-period in an OLS model specified in equation (1). The results are reported in Table 3. In column (1), we look at the number of uploaded images per month and quantify the average treatment effect as 1 less upload per month, which translates into a reduction of about 37%. In column (2) we add user-fixed effects to capture any user-specific time-invariant variation. The results remain similar and statistically indistinguishable. Further, in columns (3) and (4), we document a decrease in the extensive margin of about 30%, i.e., when we use a dependent variable that indicates whether a user has uploaded at least one image in a given month. Finally, in columns (5) and (6), we get similar results when we apply a log transformation to the dependent variable.
4.1.3 Extrapolation to aggregate effects
We can now use the average treatment effect estimates to extrapolate aggregate and long-run effects to effectively simulate counterfactuals by varying the proportion of treated users. To do this, we define three key variables: $\lambda$, the share of users in the LITE dataset; $\hat{\delta}$, the effect of inclusion in LITE; and $g$, the factual post-period growth rate in images per user. The counterfactual growth rate, had LITE not been released, is $\bar{g}=g/(1-\lambda+\lambda\hat{\delta})$. Using estimates from column (2) of Table 3, we calculate $\hat{\delta}=1-(1.1099/2.9038)=0.6$. When the LITE dataset was created in June 2020, Unsplash had a stock of $S_{0}=1,931,324$ images uploaded by $U_{0}=198,792$ users. The LITE dataset encompasses $U_{L}=8,298$ users, such that the share of users in the LITE dataset was $\lambda=U_{L}/U_{0}=0.042$. About three years later, in May 2023, Unsplash had $S_{t}=4,960,120$ images and $U_{t}=338,635$ users. Hence, the factual growth rate $g=\frac{S_{t}}{U_{t}}/\frac{S_{0}}{U_{0}}-1=14.65/9.72-1=0.51$ and therefore the counterfactual growth rate $\bar{g}=(0.51)/(1-0.042+0.042\times 0.6)=0.52$. In absolute terms, without the release of the LITE dataset, users would have uploaded 19,313 more images to Unsplash ($=S_{0}(0.52-0.51)$). We can now extrapolate that a twice-as-large LITE dataset would have reduced the flow of data by 2% (=1-0.50/0.51) compared to the factual while making the entire catalog available for commercial AI research would have reduced the flow of data by 39% (=1-0.31/0.51), corresponding to 386,264 fewer images uploaded over a three-year period.
4.1.4 Robustness
Probability weighting
Table A.2 provides a version of the baseline specification (with month and user fixed effects) but additionally weights observations inversely by the number of images the user uploaded to the platform before the LITE dataset was created. This accounts for a mechanically higher probability of having images included in the LITE dataset if a user is more prolific. The results remain very similar and statistically indistinguishable from the baseline results (columns (2), (4), and (6) in Table 3) in the sense that 90% confidence bands of the coefficients overlap.
Alternative control group definitions
Table A.3 uses the same specification as in our baseline estimations in Table 3 but with different definitions of the control group. In columns (1)–(3), we report results using a control group composed of users that did not have nature-themed (i.e., not with the keyword “nature”) and no curated images at the time when the LITE dataset was created. The effect size for the number of uploads is 47% in column (1), 33% in terms of the extensive margin in column (2), and 19% when using a specification with the log-transformed number of uploads in column (3). In columns (4)–(6), we use a control group comprised of all users whose images were not included in the LITE dataset. The effect sizes are 50%, 40%, and 21%, respectively. Note that the estimates reported in Table A.3 are not significantly different from those reported in Table 3 in that 90% confidence bands overlap.
Comparison to upload behavior on Instagram
Figure 2: Number of uploads, Instagram vs. Unsplash
A: Levels
B: Insta-Unsplash diff within control
C: Insta-Unsplash diff within treatment
Note: In Panel A, we depict the demeaned average number of uploaded images for users in the treatment group (solid) and users in the control group (dashed). Black lines indicate uploads to Unsplash, whereas grey lines indicate uploads to Instagram. We use the group-specific pre-period mean number of uploaded images to demean (2.00 [0.97] for the treatment group on Unsplash [Instagram] and 0.81 [0.90] for the control group on Unsplash [Instagram]). In Panels B and C, we plot OLS estimates of the $\delta_{\tau}$ coefficients obtained from estimating $Y_{ijt}=\sum_{\tau\in T}\nu_{\tau}+\sum_{\tau\in T}\delta_{\tau}\left(\gamma_{\tau}\times Instagram_{i}\right)+\mu_{i}+\varepsilon_{ijt}$, where $j$ indicates the treatment/control group with the number of uploads of user $i$ per month $t$ as the dependent variable. In Panel B (C), we report results from estimating this model on the subset of users in the Unsplash control group (treatment group). The dots reflect month-specific point estimates comparing upload behavior on Instagram to upload behavior on Unsplash. Standard errors are clustered at the user level, and bars indicate 90% confidence bands. In all panels, the vertical line indicates the creation of LITE 1.0.0.
One could speculate that the Unsplash research program might influence overall upload behavior on Unsplash, not only among users whose images appear in the LITE dataset (available for commercial applications) but also among those whose images are only included in the FULL dataset (available only for noncommercial research). To test this, we leverage the fact that a subset of Unsplash users is also active on the image-sharing platform Instagram. This allows us to compare the upload behavior of the same user (either with images in the LITE dataset or not) on Unsplash to their upload behavior on Instagram, which is not affected by Unsplash’s research program. Especially for professional users, Unsplash and Instagram serve as comparable platforms for building an audience, and both require contributors to grant the platform a similar license when publishing images. [^21] According to Instagram’s terms of service, users retain ownership of any intellectual property rights that they hold in the content. However, by posting to Instagram, users grant the platform a non-exclusive, royalty-free, transferable, sublicensable, worldwide license to use the content they post. Unsplash operates in a similar, yet more permissive fashion. When a user uploads an image to Unsplash, they grant anyone a license to use their image for free, including for commercial purposes, without requiring permission from or providing attribution to the user. Instagram’s terms are not explicitly clear about commercial use by third parties unless it is sublicensed, while Unsplash explicitly allows commercial use by anyone. See https://help.instagram.com/478745558852511/ and https://unsplash.com/license. For a subset of 1,084 users who disclosed this information in their Unsplash profiles, we have their Instagram account names and have gathered data on their upload history on Instagram.222222 Instagram account information is not available in the LITE or FULL datasets. We, therefore, do not have time-varying information, but we were able to get a snapshot on February 5, 2024, from which we observe Instagram accounts and then collected information about their Instagram activities. Further, technical restrictions require us to limit ourselves to the 20 most recent Instagram posts of a user. We exclude users for whom we only observe posts in the after period as we do not know whether this is because the account did not exist in the before period, whether there were no posts in the before period, or whether the user is so active that we can only see posts from the after period in our snapshot of the 20 most recent posts. Panel A of Figure 2 shows the demeaned average number of monthly uploads to Unsplash (black) and Instagram (grey). We distinguish between uploads of users who have images in the LITE dataset (“treated”, depicted with solid black and solid grey lines) and uploads of users in our preferred control group (i.e. similar types of images but they are only part of the FULL data, depicted with dashed black and grey lines). The plot suggests that the upload behavior of the control group does not differ between Instagram and Unsplash. This is confirmed in Panel B, which shows insignificant coefficients from a regression that tests, within users that are part of the Unsplash control group, whether the monthly number of uploads differs between Instagram and Unsplash.
OLS versus non-linear models
Given that our outcome variable is both a count, i.e., a non-negative integer, and zero in 86.6% of the observations, one might be worried that OLS is not the optimal choice of estimator (Wooldridge, 2023). Table A.5 shows that our baseline results are broadly robust to using a Poisson pseudo-maximum likelihood regression (PPML) approach. [^23] We do not report results from a negative binomial model because PPML is preferred for its robustness (Blackburn, 2015). However, the results indicate that PPML is sensitive to large values of the DV, which is of economic interest in our setting (i.e., prolific contributors). In addition, the interaction effect in non-linear models cannot be directly interpreted as the average treatment effect on the treated (Puhani, 2012) while simultaneously imposing additional restrictions for parallel trends to hold (Roth and Sant’Anna, 2023). For these reasons, and simplicity in the interpretation of effect sizes, we report estimates from OLS regressions throughout.
4.2 Mechanisms
In this section, we examine the potential underlying mechanisms for the baseline results. First, we provide anecdotal evidence highlighting concerns from the photography community, which we use to inform a heterogeneity analysis of the observational data. In particular, we analyze the economic motives of contributors, assessing whether professional and prolific contributors respond differently. We also examine whether the degree of exposure to the LITE dataset and growing awareness of the potential of genAI models influences user behavior. Finally, we investigate whether affected users alter the type of images they upload and whether they also change their behavior on another image-sharing platform or disengage from the profession entirely.
4.2.1 Anecdotal evidence
Discussions in online forums and blog posts offer insights into the reasons for contributors’ negative reactions to the Unsplash research program. Comments reflect a polarized response, with strong opinions on its impact on the photography industry. The predominant sentiment is negative, particularly among professional photographers who argue that Unsplash devalues their work and undermines the sustainability of their profession. Many describe the platform’s research program as exploitative, asserting that it effectively makes photographers give away their work for “exposure” while Unsplash may profit from partnerships with major companies. Others highlight how the program will contribute to the decline of traditional stock photography, making it increasingly difficult for photographers to earn a living through image licensing. Some commenters provide neutral, clarifying remarks about the dataset’s content by correcting misconceptions regarding its licensing terms and explaining that it contains metadata rather than actual image files. A minority of users express support for Unsplash, viewing it as a valuable resource for creative and research purposes, particularly for those who do not rely on photography for income. The following quotes illustrate the overall sentiment:
“If food could be obtained for free, the supermarkets would not be in business too long. Why would photographers regularly give away their licenses and claims to exclusivity to an outfit that will just parcel out their products for free? […]” (Eric on Petapixel.com)
“[…] unsplash perhpas isnt for professionals, it is for hobbiest, and semipro people who need to showcase their work. for me free for all is great way to publish my work and im just a semihobbiest ;)” (survivor303 in response to Eric on Petapixel.com)
“ […] This [Unsplash’s release of datasets] isn’t part of the evolution of digital photography as some have said. It’s just predatory. […]” (Lahmajunior on DPReview.com)
“This [Unsplash’s release of datasets] is just the normal progression of a commoditized profession… Sad, but it is the reality of photography. It is a vicious cycle, when photographers cannot make money, they don’t buy new gear. […]” (KC on Petapixel.com)
“[…] professional photography will or is already going the same path as countless other professions went before in history. Namely that the know-how of professional is getting somehow devalued by new technology, which allows for cheap mass production of a good instead of labour intensive craftsmanship. […]” (Toni Shalmonelli on DPReview.com)
We distill two key insights: First, the primarily negative reactions to Unsplash’s research program by the group of users that contribute images to the platform (i.e., photographers) seem very much in line with our baseline results. Second, the response may not be uniform but may be stronger among contributors who have more to lose from their work being freely available for AI training purposes (i.e., professionals).
4.2.2 Economic motives
Building on anecdotal evidence, we systematically test whether user behavior on Unsplash differs based on economic motives. First, we distinguish between professional and amateur contributors. The results in column (1) of Table 4 show that treated users who upload images taken with professional gear react even more negatively than others. Similarly, we find in column (2) that treated users with a badge on their profile page that says “Available for hire” reduced their uploads disproportionately compared to users without a badge. It is possible that users with professional gear and those who signal that they are available for hire make a living as photographers and are, therefore, perhaps more directly affected by the potential displacement effects of AI. Hence, monetary motives may be one of the drivers for users to contribute less to the platform after their works have been made available for AI research.
@@ 4.2.2 Economic motives
Note: The dependent variable is the number of uploads in a given month. Treated indicates whether a user was part of the LITE dataset. Pro-Gear indicates whether the user has uploaded at least one photo taken with professional gear (see section 3.2 for a detailed definition). ForHire indicates whether a user has chosen to show “Available for hire” on their Unsplash profile page. The number of observations differs from those in Table 3 because we do not observe Pro-Gear for all images and ForHire for all users. Fixed effects for upload month of the image and user account. Standard errors are clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$ $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
Further, it is worth noting that we observe that users in the treatment group are three times more likely to be contributors in the Unsplash+ program (t-statistic in a two-sided t-test is 6.13). This program enables users to gain monetary compensation and explicitly disallows the use of their works for AI training. [^24] See https://unsplash.com/plus/license and https://unsplash.com/blog/contribute-to-unsplash/. Hence, we interpret the substantially higher uptake of Unsplash+ among users who had at least one image in the LITE dataset as a response to counteract economic forces and to avoid the utilization of their works in AI applications.
@@
Another, possibly correlated dimension of user heterogeneity concerns upload activity, which is an important metric for gaining insights into platform governance. In Table A.6, we show that the main effect reported in column (1) of Table 3 varies by percentile of a user’s total upload activity. The point estimates in columns (1)–(4) of Table A.6 are smaller and not always significantly different from the estimate in column (1) of Table 3 in the sense that 90% confidence bands overlap. However, in absolute terms, the implied effect sizes are larger (column 2: -45%, column 3: -44%, and column 4: -40%) compared to the -30% of the baseline. Given that we observe the universe of activity on Unsplash, we prefer not to remove particularly prolific users and report estimates of average treatment effects based on all users as our more conservative baseline results.
4.2.3 Treatment intensity
Table 5: Results: Changes in uploads, treatment intensity
Uploads
(1) (2)
Post $\times$ TreatedSingle -0.3059∗∗∗
Post $\times$ TreatedMultiple -2.2830∗∗∗
Post $\times$ Treated -1.0056∗∗∗
PostAug22 $\times$ Treated -0.4054∗∗∗
Mean Treated Before 2.9038
Mean TreatedSingle Before 1.4011
Mean TreatedMultiple Before 5.0982
Month FE Yes Yes
User FE Yes Yes
Observations 612,662 612,662
Note: The dependent variable is the total number of uploads per month. Post indicates the entire period after the creation of LITE. TreatedSingle indicates whether a user had exactly one image in the LITE dataset. TreatedMultiple indicates whether a user had more than one image in the LITE dataset. Treated indicates whether a user was part of the LITE dataset. PostAug22 indicates the period after August 2022, when groundbreaking genAI for images was released. Fixed effects for upload month of the image and user account. Standard errors are clustered at the user level in parentheses. $p<0.10$, ∗∗ $p<0.05$∗∗ $p<0.01$
The results in Table 5 provide more evidence of the potential mechanisms through which the inclusion in the LITE dataset induces users to reduce their upload activity.
In column (1), we distinguish between users that were affected more or less intensively by the release of the LITE dataset. In particular, we distinguish between users who had one image included (about 60%) and users who had multiple images included (about 40%). Conditional on more than one image in the LITE dataset, the average number of images per user is 5.96 (standard deviation 14.18 and maximum 419). The estimated effect size is 22% for users with only one image included versus 45% for users with multiple images.. This difference is statistically significant. These results provide further evidence that users must have been aware of their images being included in the LITE dataset because otherwise, we would not expect a stronger response from more heavily treated users.
In column (2), informed by Figure 1, we split the post-period into two parts. The first 24 months correspond to the period between June 2020 and August 2022, which garnered modest public interest in AI and AI art in particular. After August 2022, however, there was a substantial increase in media coverage of the topic, e.g., a New York Times article titled “We Need to Talk About How Good A.I. Is Getting”, which may have further raised awareness about genAI in the Unsplash community. [^25] See https://www.nytimes.com/2022/08/24/technology/ai-technology-progress.html The timing also coincides with the public release of Stable Diffusion, the first text-to-image generator that could be run on consumer-grade machines, democratizing access to genAI technology after similar technology released a few months earlier had only restricted public access (e.g., DALL-E and Midjourney). [^26] See https://stability.ai/news/stable-diffusion-announcement. We find a treatment effect of 34% before August 2022 and 49% thereafter, which corresponds to a significant 40% increase in the treatment effect. This suggests that, as awareness of the commercial potential of image generation AI rises, photographers reduce their contributions to Unsplash and, therefore, to the flow of AI training data.
4.2.4 Differentiation or exit?